Please help and show all work on how to get answer. Appreciate it!!
4. A 10:1 step up transformer is connected to a 60 volt generator as shown. The resistance in the secondary circuit is 100 ohms. Find: a. The voltage in the secondary circuit. [3] b. The current in the secondary circuit. [3 c. The current in the primary circuit. [3]
Answer #1

a)   voltage in secondary circuit = 60 * 10 = 600 V

b)    current in secondary circuit = 600/100 = 6 A

c)   current in primary circuit =   6 * 10 = 60 A